OpenStudy (anonymous):

The Lewiston Landscaping Company charged Mr. Rabinowitz $100 for mowing his rectangular lawn, which is 20 feet long and 12 feet wide. If Mr. Nabors's lawn is also rectangular and measures 15 feet long and 10 feet wide, and the price charged for mowing a lawn is proportional to its area, which of the following expressions can be used to calculate how much (x) the company will charge Mr. Nabors?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (littlebird):

What are the expressions?

OpenStudy (anonymous):

1.100/x=32/25 2.100/150=x/240 3.180/200=150/x 4.100/240=x/150 5.240/150=x/150

OpenStudy (anonymous):

?

OpenStudy (anonymous):

very helpful thanks